12. 20 mL of a 10% solution was diluted to 100 mL in a 100 mL volumetric flask. What is the concentration of the new solution?
1
Expert's answer
2020-04-22T09:20:16-0400

w2 = V1×w1/V2 = 20×10%/100 = 2%.